Write an equation for the circle with center (-4,√7) and radius of 5 units.
Thank You

Answer by fractalier(2101)

The general form is
(x - h)^2 + (y - k)^2 = r^2
so plugging in we have
(x + 4)^2 + (y - √7)^2 = 25